### v2

# Examples of Ricci limit spaces with infinite holes

Abstract.
Let $n\geq 3$, $\lambda\in\mathbb{R}$, and $(X,h)$ be an $n$-dimensional smooth complete Riemannian manifold with $\mathrm{Ric}_{h}>\lambda$.
In this paper, we construct, for each given $\epsilon>0$, a sequence of $(n+2)$-dimensional manifolds $(M_{i},g_{i})\stackrel{{\scriptstyle GH}}{{\longrightarrow}}(X_{\epsilon},d_{\epsilon})$ with $\mathrm{Ric}_{g_{i}}>\lambda$, such that $d_{GH}(X,X_{\epsilon})\leq\epsilon$, and $X_{\epsilon}$ is homeomorphic to the space obtained by removing an infinite number of balls from $X$. Hence $X_{\epsilon}$ has dense boundary with an infinite number of connected components. Moreover, $X_{\epsilon}$ has no open subset which is topologically a manifold. This generalizes Hupp-Naber-Wang’s result [11] from $4$-dimensional case to the case of dimension $n\geq 3$.
Our construction differs from that of Hupp-Naber-Wang. In their approach, Hupp-Naber-Wang considered doing an infinite number of blow-ups on the local complex surface structure of $X$, thus relying on the $4$-dimensional condition. However, our approach avoids the use of local complex surface structure of $X$, allowing us to construct in the general case of dimensions greater than or equal to $3$. As a corollary, we provide a solution to an open problem posed by Naber [16, Open Problem 3.4] in the $3$-dimensional case.
1. Introduction
In the theory of Gromov-Hausdorff limits, a significant focus lies in describing the geometry and topology of measured Gromov-Hausdorff limit spaces as following:
(1) $(M_{i},g_{i},\nu_{i},p_{i})\stackrel{{\scriptstyle GH}}{{\longrightarrow}}(X^{n},d,\nu,p),\;\;\mathrm{Ric}_{g_{i}}\geq-\lambda,\;\;\nu_{i}=\frac{1}{\mathrm{Vol}(B_{1}(p_{i}))}d\nu_{g_{i}},$
where $n\in\mathbb{N}$ is the rectifiable dimension of $(X,d,\nu)$, and all $M_{i}$ have the same dimension. Note that Colding-Naber [6] proved that the limit $(X,d,\nu)$ is $n$-rectifiable for a unique $n$.
In the non-collapsed case, which is to say $\mathrm{Vol}(B_{1}(p_{i}))\geq v>0$, the works of Colding [5, Theorem 0.1] and Cheeger-Colding [2, Theorem 5.9] show that the measure $\nu=c\mathcal{H}^{n}$ for some $c>0$, and the set of $n$-regular points $R_{n}(X)\subset X$ is a subset of full-$\nu$ measure, where $\mathcal{H}^{n}$ is the $n$-dimensional Hausdorff measure on $(X^{n},d)$, and a point $x\in X$ is called a $k$-regular point if and only if every tangent cone of $X$ at $x$ is isometric to $\mathbb{R}^{k}$. Then a Reifenberg type result of Cheeger-Colding [2, Theorem 5.14] shows that the limit space has a manifold structure away from a codimension $2$ set, which is the complement of an open neighborhood of $R_{n}(X)$. More recently, the work of Cheeger-Jiang-Naber [4, Theorem 1.14] was able to prove that in this case the codimension $2$ Hausdorff measure of the subset we removed can be finite.
In the collapsed case, which is to say $\lim\limits_{i\to\infty}\mathrm{Vol}(B_{1}(p_{i}))=0$, Colding-Naber [6] proved that the set of $n$-regular points $R_{n}(X)\subset X$ is a subset of full-$\nu$ measure. However, Pan-Wei’s example [17] has shown that the Hausdorff dimension of $X\backslash R_{n}(X)$ can be greater than the Hausdorff dimension of $R_{n}(X)$. By using their almost $n$-splitting map, Cheeger-Colding [3, Theorem 3.26] proved that there exists a subset $U$ of $R_{n}(X)$ such that $\nu(U)>0$ and $U$ is bi-Lipschitz to a subset of $\mathbb{R}^{n}$. However, in general, $U$ is not an open subset of $X$.
A natural question arises: if $x\in R_{n}(X)$, does there exist an open neighborhood of $x$ homeomorphic to an open subset of $\mathbb{R}^{n}$, similar to the non-collapsed case? This leads to the following famous problem:
Problem 1.1 ([16, Open Problem 3.4]).
Let $(M_{i},g_{i},\nu_{i},p_{i})\stackrel{{\scriptstyle GH}}{{\longrightarrow}}(X^{n},d,\nu,p)$ with $\mathrm{Ric}_{g_{i}}\geq-\lambda$ and $\mathrm{Vol}(B_{1}(p_{i}))\to 0$, then is there an open subset of full-$\nu$ measure $R(X)\subset X$ which is homeomorphic to an $n$-manifold?
It’s worth noting that for $n=1$, the answer of Problem 1.1 is affirmative [10, 7].
Recently, Hupp-Naber-Wang [11] demonstrated that for any $4$-dimensional manifold with Ricci lower bound, there exists a sequence of $6$-dimensional manifolds with a uniform Ricci lower bound such that the Gromov-Hausdorff limit is close to the given $4$-dimensional manifold, but the limit space has no open subset which is topologically a manifold. In topological sense, the Ricci limit space they constructed is the blowing up of a dense subset in the given $4$-dimensional manifold. As a corollary, by considering the product of a $4$-dimensional manifold and an $(n-4)$-dimensional manifold, they provided a negative answer of Problem 1.1 when $n\geq 4$.
However, Problem 1.1 is still open when $n=2$ or $3$.
On one hand, the blow-up surgery used by Hupp-Naber-Wang requires local complex surface structures and cannot be applied to $2$ or $3$-manifolds. On the other hand, Simon-Topping’s result [19], about the manifold structure of $3$-dimensional non-collapsing Ricci limit spaces indicates that there may be some additional topological regularity in the $3$-dimensional case compared to higher dimensions. Here we pay special attention to the $3$-dimensional case, because in many cases people tend to focus on the geometry of $6$-manifolds collapse to a $3$-space, such as the famous metric Strominger-Yau-Zaslow conjecture [13].
In this paper, we construct a sequence of $(n+2)$-dimensional manifolds for any given $n$-dimensional manifolds $(n\geq 3)$ with Ricci lower bound such that the limit space has no open subset which is a topological manifold. The following theorem is our main theorem in this paper.
Theorem 1.1.
Let $(X,h)$ be an $n$-dimensional smooth complete boundary free Riemannian manifold with $n\geq 3$ and $\mathrm{Ric}_{h}>\lambda\in\mathbb{R}$. Then for each $\epsilon>0$, there exists a metric space $(X_{\epsilon},d_{\epsilon})$ such that
  1. (a).
    $d_{GH}(X,X_{\epsilon})\leq\epsilon$,
  2. (b).
    $(X_{\epsilon},d_{\epsilon})$ is $n$-rectifiable,
  3. (c).
    There exists a sequence of $(n+2)$-dimensional boundary free complete manifolds $(M_{i},g_{i})\stackrel{{\scriptstyle GH}}{{\longrightarrow}}(X_{\epsilon},d_{\epsilon})$ with $\mathrm{Ric}_{g_{i}}>\lambda$,
  4. (d).
    For any open subset $U$ of $X_{\epsilon}$, the homology group $H_{n-1}(U;\mathbb{Z})$ is infinitely generated. Consequently, every open subset of $X_{\epsilon}$ is not a topological manifold.
Topologically, the main difference between our construction and that of Hupp-Naber-Wang is that we use the plumbing gluing surgery to puncture holes on the manifold the given manifold $X$, instead using the blow-up surgery on the local complex structure of $X$. By using this surgery and careful analysis near the boundaries of the balls we removed, one can construct for the general case of dimension $n\geq 3$. Thus for manifolds of dimension greater or equal to $5$, our method can also construct Ricci limit spaces that approximate any given manifold with dense homology, without necessitating a product structure.
Remark.
Now we can refine Hupp-Naber-Wang’s answer to Problem 1.1. Specifically, the following facts are now known:
  • When $n=1$, the Ricci limit space of rectifiable dimension $n$ is always a topological manifold with boundary. For more details, see Honda [10] or Chen [10].
  • When $n\geq 3$, Theorem 1.1 implies that one can approximate any $n$-dimensional Riemannian manifold with lower Ricci bound by $n$-rectifiable Ricci limit spaces with bad topology.
Now we consider the boundary of Ricci limit spaces. Let $(X^{n},d,\nu,p)$ be a Ricci limit space. Then for $10\leq k\leq n-$ 1, one can define
(2) $\mathcal{S}^{k}=\bigg{\{}x\in X:\textrm{ no tangent cone at $x$ splits off $\mathbb{R}^{k+1}$ }\bigg{\}}.$
Now we can define the reduced boundary of $X$ by
(3) $\partial^{*}X=\mathcal{S}^{n-1}\backslash\mathcal{S}^{n-2}.$
There are two definitions of boundary. The definition of boundary given by De Philippis-Gigli [8], $\partial_{\rm DPG}X$, is just the closure of the reduced boundary. The other definition of a boundary, provided by Kapovitch-Mondino [12], $\partial_{\rm KM}X$, is defined by induction on tangent cones. See also [1] for more details about boundaries.
The motivation for studying the geometry of boundaries comes from the research about $\mathrm{RCD}$ spaces. Note that Cheeger-Colding [2, Theorem 6.1] proved that the boundary of non-collapsed Gromov-Hausdorff limits of boundary free manifolds is always empty. As a corollary of Theorem 1.1, our construction also gives examples where boundaries are both dense and contain an infinite number of connected components.
Corollary 1.1.
Let $n\geq 3$. Then there exists a sequence of $(n+2)$-dimensional boundary free complete manifolds $(M_{i},g_{i})\stackrel{{\scriptstyle GH}}{{\longrightarrow}}(X^{n}_{\infty},d_{\infty})$ with $\mathrm{Ric}_{g_{i}}\geq-1$, such that $\partial_{KM}X_{\infty}$ is both dense and contains an infinite number of connected components. Moreover, $\partial_{\rm DPG}X_{\infty}=X_{\infty}$.
Remark.
We thank Professor Aaron Naber for pointing out this interesting corollary.
This paper is organized as follows. At first, we will give an outline of our construction in Section 2. In Section 3, we collect some preliminary results. Then we will construct the local model in Section 4. The construction of $(M_{i},g_{i})$ is presented in Section 5. Finally, the proofs of Theorem 1.1 and Corollary 1.1 are contained in Section 6.
Acknowledgement. The author wants to express his deep gratitude to Professor Gang Tian for constant encouragement. The author would also be grateful to Professor Aaron Naber, Professor Wenshuai Jiang and Professor Jianchun Chu for valuable conversations. Additionally, the author thanks Professor Xiaochun Rong and Professor Guofang Wei for their interests in this note. Furthermore, the author expresses gratitude to Professor Xin Fu and Professor Jiyuan Han for their warm hospitality during his visit to Westlake University. In particular, the author would like to thank Daniele Semola for pointing out some errors in Introduction of the first edition.
2. Outline of the construction
At first, we describe our construction in the topological sense. For the moment, let’s set aside considerations of Ricci curvature and smoothness.
Our approach to puncture holes in the manifold $X$ is to consider plumbing gluing surgery on higher dimensional manifolds. Although the manifolds in the statement of Theorem 1.1 are manifolds without boundary, the new boundaries will appear after we puncture holes on the manifold $X$. Thus we assume here that $X$ is a manifold with boundary in the following description.
Let $n,k\geq 1$, $X$ be an $n$-dimensional manifold with boundary, $M$ be an $(n+k-1)$-dimensional manifold without boundary, and $F:M\to X$ be a continuous map such that:
  • The restriction of $F$ on $F^{-1}(\mathrm{int}(X))\subset M$ gives a ${\mathbb{S}}^{k-1}$-bundle structure on $F^{-1}(\mathrm{int}(X))$.
  • The pre-image $F^{-1}(x)$ is a single point if $x\in\partial X$.
Choosing a pre-compact open subset $U$ of $\mathrm{int}(X)$ such that $\bar{U}$ is homeomorphic to $\bar{B}^{n}_{1}(0)\subset\mathbb{R}^{n}$. Then $\partial F^{-1}(\bar{U})$ is homeomorphic to ${\mathbb{S}}^{n-1}\times{\mathbb{S}}^{k-1}$. Clearly, $\partial\left({\mathbb{S}}^{n-1}\times\bar{B}^{k}_{1}(0)\right)$ is also homeomorphic to ${\mathbb{S}}^{n-1}\times{\mathbb{S}}^{k-1}$. Hence we can get a new $(n+k-1)$-dimensional manifold $M^{\prime}$ by gluing ${\mathbb{S}}^{n-1}\times\bar{B}^{k}_{1}(0)$ to $M\backslash F^{-1}(U)$.
Now we consider the Euclidean distance function $d(y)=d_{\rm Euc}(y,0)$ on $\bar{B}^{k}_{1}(0)$. Clearly, $d:\bar{B}^{k}_{1}(0)\to[0,1]$ gives a ${\mathbb{S}}^{k-1}$-bundle structure on $d^{-1}((0,1])=\bar{B}^{k}_{1}(0)\backslash\{0\}$, and $d^{-1}(0)$ is a single point $\{0\}$. Then we can get a map $\pi_{d}:{\mathbb{S}}^{n-1}\times\bar{B}^{k}_{1}(0)\to{\mathbb{S}}^{n-1}\times[0,1]$ such that $\pi_{d}^{-1}$ gives a ${\mathbb{S}}^{k-1}$-bundle structure on $\pi_{d}^{-1}({\mathbb{S}}^{n-1}\times(0,1])={\mathbb{S}}^{n-1}\times\left(\bar{B}^{k}_{1}(0)\backslash\{0\}\right)$, and $\pi_{d}^{-1}(y)$ is a single point if $y\in{\mathbb{S}}^{n-1}\times\{0\}$. Then we can obtain an $n$-dimensional manifold $X^{\prime}$ with boundary $\partial X^{\prime}=\partial X\bigsqcup{\mathbb{S}}^{n-1}$ by gluing ${\mathbb{S}}^{n-1}\times[0,1]$ to $X\backslash U$, and get a continuous map $F^{\prime}:M^{\prime}\to X^{\prime}$ by gluing $F$ and $\pi_{d}$.
Actually, the metric on $\bar{B}^{k}_{1}(0)$ we used is isometric to the cone metric on $C({\mathbb{S}}^{k-1}_{\theta})$ for some $\theta\in(0,1)$. We then shrink the restriction of the metric on the ${\mathbb{S}}^{k-1}$-fibers. As the constant $\theta$ decreases, $C({\mathbb{S}}^{k-1}_{\theta})$ converges to the interval $[0,1]$.
It is easy to see that the data $(X^{\prime},M^{\prime},F^{\prime})$ also satisfies the conditions above, allowing us to replace $(X,M,F)$ by $(X^{\prime},M^{\prime},F^{\prime})$. By repeating this process, we can puncture an infinite number of holes in $X$, and the topology of the Ricci limit space that we have constructed is the space obtained by removing an infinite number of balls from $X$.
To make our construction easier to understand, we give here a graph for the case where $n=2$, $k=1$, $X_{0}=\bar{\mathbb{D}}_{1}(0)$, $M_{0}={\mathbb{S}}^{2}$ and $F_{0}$ is the standard projection. Here, $M_{i}$ is getting squashed, causing the Gromov-Hausdorff distance between $M_{i}$ and $X_{i}$ to decrease.
[…]
The above graph is to provide some intuition. Now we give an outline of our construction.
  1. (I).
    Write $(X_{0},h_{0})=(X,h)$, $(M_{0},g_{0})=\left(X\times{\mathbb{S}}^{2},h+(1+|\lambda|)^{-10}g_{{\mathbb{S}}^{2}}\right)$, $f_{0}=(1+|\lambda|)^{-5}$, and $W_{0}=\emptyset\subset M_{0}$. Then we have $\mathrm{Ric}_{g_{0}}>\lambda$.
  2. (II).
    Let $n\geq 3$, $(X_{i},h_{i})$ be an $n$-dimensional smooth Riemannian manifold with boundary, $f_{i}$ be a non-negative function on $X_{i}$, $f_{i}>0$ on $\mathrm{int}(X_{i})$, $(M_{i},g_{i})$ be an $(n+2)$-dimensional smooth Riemannian manifold without boundary, $\mathrm{Ric}_{g_{i}}>\lambda$, and $W_{i}\subset M_{i}$ be an $(n-1)$-dimensional submanifold such that
    $(M_{i}\backslash W_{i},g_{i})\cong(\mathrm{int}(X_{i})\times{\mathbb{S}}^{2},h_{i}+f_{i}^{2}g_{{\mathbb{S}}^{2}}).$
  3. (III).
    Let $U_{i}$ be a small pre-compact open subset of $\mathrm{int}(X_{i})$, such that $\bar{U}_{i}$ diffeomorphic to $\bar{B}^{n}_{1}(0)$.
  4. (IV).
    Construct a metric $h_{i+1}$ and a non-negative function $f_{i+1}$ on $X_{i}\backslash U_{i}$, such that $f_{i+1}=0$ on $\partial U_{i}$, and the completion of $((\mathrm{int}(X_{i})\backslash U_{i})\times{\mathbb{S}}^{2},h_{i+1}+f_{i+1}^{2}g_{{\mathbb{S}}^{2}})$ near $\partial U_{i}$ is a smooth Riemannian manifold without boundary. Moreover, we can assume that $h_{i+1}=h_{i}$ and $f_{i+1}=\delta f_{i}$ outside a small neighborhood of $\partial U_{i}$.
  5. (V).
    Consider the completion $(Y,d)$ of $((\mathrm{int}(X_{i})\backslash U_{i})\times{\mathbb{S}}^{2},h_{i+1}+f_{i+1}^{2}g_{{\mathbb{S}}^{2}})$. Then there is a natural embedding $W_{i}\to Y$, and tangent cone at any point in $W_{i}$ is always isometric to $C({\mathbb{S}}^{2}_{\delta})\times\mathbb{R}^{n-1}$.
  6. (VI).
    Modify $f_{i+1}$ near $W_{i}$ such that the completion of $((\mathrm{int}(X_{i})\backslash U_{i})\times{\mathbb{S}}^{2},h_{i+1}+f_{i+1}^{2}g_{{\mathbb{S}}^{2}})$ near $\partial W_{i}$ becomes a smooth Riemannian manifold without boundary. See also Corollary 4.2.
  7. (VII).
    Let $(M_{i+1},g_{i+1})$ be the completion of $((\mathrm{int}(X_{i})\backslash U_{i})\times{\mathbb{S}}^{2},h_{i+1}+f_{i+1}^{2}g_{{\mathbb{S}}^{2}})$, and $(X_{i+1},h_{i+1},f_{i+1})=(X_{i}\backslash U_{i},h_{i},f_{i})$. Then we do the process (ii)-(v) again and again.
  8. (VIII).
    Since $d_{GH}(M_{i},X_{i})=\Psi(i^{-1})$, we see that the Gromov-Hausdorff limits $\lim M_{i}=\lim X_{i}=X_{\epsilon}$. By composing the inclusion maps $X_{i+1}=X_{i}\backslash U_{i}\to X_{i}$, we can obtain a Lipschitz map $X_{\epsilon}\to X_{i}$. Then we can prove Theorem 1.1.
Remark.
From the above construction, we see that by replacing $\mathbb{S}^{2}$ with $\mathbb{S}^{m}$, $m\geq 2$, one can obtain in the similar way a sequence of $(n+m)$-dimensional manifolds $(M_{i},g_{i})\stackrel{{\scriptstyle GH}}{{\longrightarrow}}(X_{\epsilon},d_{\epsilon})$ with $\mathrm{Ric}_{g_{i}}>\lambda$ and $X_{\epsilon}$ has infinite holes.
3. Preliminary
3.1. Double warped products
In this section, we recall the Ricci curvature of double warped products.
Let $\varphi,\phi$ be smooth nonnegative functions on $[0,\infty)$ such that $\varphi,\phi$ are positive on $(0,\infty)$,
(4) $\phi(0)>0,\;\phi^{\mathrm{(odd)}}(0)=0,$
and
(5) $\varphi(0)=0,\;\varphi^{\prime}(0)=1,\;\varphi^{\mathrm{(even)}}(0)=0.$
Then we can define a Riemannian metric on $\mathbb{R}^{3}\times{\mathbb{S}}^{n}$ by
$g_{\varphi,\phi}(r)=dr^{2}+\varphi(r)^{2}g_{{\mathbb{S}}^{2}}+\phi(r)^{2}g_{{\mathbb{S}}^{n}}.$
See also [18, Proposition 1.4.7] for more details.
Let $X_{0}=\frac{\partial}{\partial r}$, $X_{1}\in T_{{\mathbb{S}}^{2}}$, and $X_{2}\in T_{{\mathbb{S}}^{n}}$ be unit vectors. Then the Ricci curvature of $g_{\varphi,\phi}$ can be expressed as following.
Lemma 3.1.
Let $\varphi,\phi$ and $g_{\varphi,\phi}$ be as above. Then the Ricci curvature tensor of $g_{\varphi,\phi}$ can be determined by
(6) $\mathrm{Ric}_{g_{\varphi,\phi}}\left(X_{0}\right)$ $=$ $-\left(2\frac{\varphi^{\prime\prime}}{\varphi}+n\frac{\phi^{\prime\prime}}{\phi}\right)X_{0},$
(7) $\mathrm{Ric}_{g_{\varphi,\phi}}\left(X_{1}\right)$ $=$ $\left[-\frac{\varphi^{\prime\prime}}{\varphi}+\frac{1-(\varphi^{\prime})^{2}}{\varphi^{2}}-n\frac{\varphi^{\prime}\phi^{\prime}}{\varphi\phi}\right]X_{1},$
(8) $\mathrm{Ric}_{g_{\varphi,\phi}}\left(X_{2}\right)$ $=$ $\left[-\frac{\phi^{\prime\prime}}{\phi}+(n-1)\frac{1-(\phi^{\prime})^{2}}{\phi^{2}}-2\frac{\varphi^{\prime}\phi^{\prime}}{\varphi\phi}\right]X_{2}.$
Proof.
One can conclude it by a straightforward calculation. See also [18, Subsection 4.2.4]. ∎
Then we have the following gluing lemma for double warped products.
Lemma 3.2.
Let $\varphi,\phi$ and $g_{\varphi,\phi}$ be as above.
Assume that $\varphi^{\prime}>0,\phi^{\prime}>0$ on the interval $(r_{0}-2\delta,r_{0}+2\delta)\subset(0,\infty)$, and $\varphi,\phi$ are smooth on $(r_{0}-2\delta,r_{0})\cup(r_{0},r_{0}+2\delta)$. Suppose that $\varphi,\phi$ have smooth limits on both sides of $r_{0}$, $\varphi^{\prime}_{+}(r_{0})\leq\varphi^{\prime}_{-}(r_{0})$, and $\phi^{\prime}_{+}(r_{0})\leq\phi^{\prime}_{-}(r_{0})$. Then for any $\epsilon>0$, we can find smooth $\tilde{\varphi}$, $\tilde{\phi}$ such that $|\tilde{\varphi}-\varphi|+|\tilde{\phi}-\phi|\leq\epsilon$ on $(r_{0}-2\delta,r_{0}+2\delta)$, and $\tilde{\varphi}=\varphi$, $\tilde{\phi}=\phi$ outside $(r_{0}-\delta,r_{0}+\delta)$, and $\inf\mathrm{Ric}_{g_{\tilde{\varphi},\tilde{\phi}}}\geq\inf\mathrm{Ric}_{g_{\varphi,\phi}}-\epsilon$.
This is a consequence of Menguy’s $C^{1}$ gluing lemma [15, Lemma 1.170]. See also [11, Lemma 3.1].
In general, if we consider the warped product $(M\times{\mathbb{S}}^{n},g=g_{M}+f^{2}g_{{\mathbb{S}}^{n}})$, then the Ricci curvature operator can be given by
(9) $\mathrm{Ric}_{g}|_{TM}$ $=$ $\mathrm{Ric}_{g_{M}}-n\frac{\mathrm{Hess}_{g}f}{f},$
(10) $\mathrm{Ric}_{g}|_{T{\mathbb{S}}^{n}}$ $=$ $\left(\frac{1}{f^{2}}-(n-1)\frac{|\nabla_{g_{M}}f|_{g}^{2}}{f^{2}}-\frac{\Delta_{g}f}{f}\right)f^{2}g_{{\mathbb{S}}^{n}}.$
As a corollary, one can conclude that $\mathrm{Ric}_{g_{M}+f^{2}g_{{\mathbb{S}}^{n}}}\geq\lambda$, if and only if for any $c\in(0,1]$, $\mathrm{Ric}_{g_{M}+c^{2}f^{2}g_{{\mathbb{S}}^{n}}}\geq\lambda$. See also [11, Remark 3.2].
3.2. Adding conical singularities
Now we recall the lemma Hupp-Naber-Wang used in [11] to add conical singularities to the given manifold.
Lemma 3.3 ([11, Lemma 5.1]).
Let $(M_{B},g_{B})$ be an $n+1$-dimensional Riemannian manifold, $n\geq 2$, $p\in M_{B}$, and $f$ be a positive smooth function on $B_{2}^{n+1}(p)$. Assume that $B_{3}^{n+1}(p)$ is pre-compact in $(M_{B},g_{B})$, and the warped product space $(B_{2}^{n+1}(p)\times{\mathbb{S}}^{2},g)$ with metric $g=g_{B}+f^{2}g_{{\mathbb{S}}^{2}}$ satisfying
(11) $\mathrm{Ric}_{g}>\lambda\,,\mathrm{inj}_{g_{B}}(p)$ $\geq 2\,,$
(12) $|\mathrm{Rm}_{g_{B}}|_{g_{B}}\,,\,|\nabla\mathrm{Rm}_{g_{B}}|_{g_{B}}\,,\,|\nabla^{2}\mathrm{Rm}_{g_{B}}|_{g_{B}}$ $\leq\frac{1}{2}\,,$
(13) $|\nabla_{g_{B}}\ln f|_{g_{B}}\,,\,|\nabla^{2}_{g_{B}}\ln f|_{g_{B}}\,,\,|\nabla^{3}_{g_{B}}\ln f|_{g_{B}}$ $\leq\frac{1}{2}\,.$
[…]
  1. (1)
    The Ricci lower bound $\mathrm{Ric}_{\hat{g}}>\lambda-C(n)\epsilon$ holds for $\hat{r}/2\leq r\leq 2$  ,
  2. (2)
    $\hat{g}=g_{B}+\hat{\delta}^{2}f^{2}g_{{\mathbb{S}}^{2}}$, $\forall r\in[1,2]$ ,
  3. (3)
    $\hat{g}=dr^{2}+(1-\epsilon)^{2}r^{2}g_{{\mathbb{S}}^{n}}+\delta^{2}r^{2\alpha}g_{{\mathbb{S}}^{2}}$, $\forall r\leq\hat{r}$ ,
  4. (4)
    The identity map $\text{Id}:(B_{2}(p),\hat{g}_{B})\to(B_{2}(p),g_{B})$ is $(1+2\epsilon)$-bi-Lipschitz.
This is a slight generalization of Hupp-Naber-Wang’s lemma. Although they chose $n=4$ in their original statement, the proof is essentially verbatim.
4. Construction of the local model spaces
In this section, we construct some local model spaces by double warped products.
At first, we construct a metric $g_{\varphi,\phi}$ on $\mathbb{R}^{3}\times{\mathbb{S}}^{n}$ such that the asymptotic cone is $C({\mathbb{S}}^{n}_{1-\epsilon})$.
Lemma 4.1.
Let $m=2$ and $n\geq 2$. Then for any $0<\epsilon\leq\frac{1}{100}$, $0<\alpha\leq\alpha_{0}(n,\epsilon)$ and $0<\delta\leq\delta_{0}(n,\epsilon,\alpha)$, we can find positive constants $R(n,\epsilon,\alpha)$, $C(n,\epsilon,\alpha)$ and functions $\varphi,\phi$ satisfying (4) and (5), such that $\mathrm{Ric}_{g_{\varphi,\phi}}>0$, $\phi(0)=1$, $\phi|_{[R,\infty)}=(1-\epsilon)(r+C)$, $\varphi(0)=0$, and $\varphi|_{[R,\infty)}=\delta(r+C)^{\alpha}$.
Moreover, there exists a small constant $\tau=\tau(n,\epsilon,\alpha)>0$ such that $\mathrm{Ric}_{g_{\varphi,\phi}}\geq\frac{\tau}{r^{2}+1}$.
Remark.
Note that $R,C$ and $\tau$ are independent of $\delta$.
Proof.
Our construction is divided into three parts.
Part 1. Let $\varphi_{0}(r)=\sin(r)$ and $\phi_{0}(r)=1+r^{4}$ on $\left[0,\frac{1}{9n}\right]$. Then Lemma 3.1 implies that
$\mathrm{Ric}_{g_{\varphi_{0},\phi_{0}}}\left(X_{0},X_{0}\right)$ $=$ $2-n\frac{12r^{2}}{1+r^{4}}\geq 2-\frac{12}{81}>0,$
$\mathrm{Ric}_{g_{\varphi_{0},\phi_{0}}}\left(X_{1},X_{1}\right)$ $=$ $\varphi_{0}^{2}\left[2-n\frac{4\cos(r)r^{3}}{(1+r^{4})\sin(r)}\right]\geq\varphi^{2}\left(2-4nr^{2}\right)>0,$
$\mathrm{Ric}_{g_{\varphi_{0},\phi_{0}}}\left(X_{2},X_{2}\right)$ $=$ $\phi_{0}^{2}\left[-\frac{12r^{2}}{1+r^{4}}+(n-1)\frac{1-16r^{6}}{(1+r^{4})^{2}}-2\frac{4\cos(r)r^{3}}{(1+r^{4})\sin(r)}\right]$
$\geq$ $\phi_{0}^{2}\left(-\frac{12}{81}+\frac{1}{2}-\frac{8}{81}\right)>0.$
Note that $\tan r\geq r$, $\forall r\in(0,1]$. Clearly, we have $\phi_{0}^{\prime}(\frac{1}{10n})=\frac{1}{250n^{3}}$ and $\varphi_{0}^{\prime}(\frac{1}{10n})=\cos(\frac{1}{10n})\geq\frac{1}{2}$.
Write $c_{1}=(10n)^{-3}$ and $R_{1}=100c^{-1}$. Then we can construct $\varphi_{1},\phi_{1}$ on $[0,\infty)$ such that
[…]
and
[…]
Then for $\frac{1}{10n}\leq r\leq\frac{1}{9n}$, we have
$\mathrm{Ric}_{g_{\varphi_{1},\phi_{1}}}\left(X_{0},X_{0}\right)$$\geq$ $\frac{2\sin(r)}{r}-12nr^{2}>0,$
$\mathrm{Ric}_{g_{\varphi_{1},\phi_{1}}}\left(X_{1},X_{1}\right)$$\geq$ $\varphi_{1}\left[\sin(r)-4nr^{3}\right]>0,$
$\mathrm{Ric}_{g_{\varphi_{1},\phi_{1}}}\left(X_{2},X_{2}\right)$ $=$ $\phi_{0}\left[-12r^{2}+(n-1)\frac{1-16r^{6}}{(1+r^{4})^{2}}-\frac{8r^{3}}{(1+r^{4})\sin(\frac{1}{10n})}\right]$
$\geq$ $\phi_{1}\left(-\frac{12}{81}+\frac{1}{2}-\frac{10}{81}\right)>0.$
Similarly, when $\frac{1}{9n}\leq r\leq R_{1}$, we have
$\mathrm{Ric}_{g_{\varphi_{1},\phi_{1}}}\left(X_{0},X_{0}\right)$ $>$ $0,$
$\mathrm{Ric}_{g_{\varphi_{1},\phi_{1}}}\left(X_{1},X_{1}\right)$ $=$ $\varphi_{1}\left[-\varphi_{1}^{\prime\prime}+\frac{1-(\varphi_{1}^{\prime})^{2}}{\varphi_{1}}-n\frac{\varphi_{1}^{\prime}\phi_{1}^{\prime}}{\phi_{1}}\right]$
$>$ $\varphi_{1}\left[\frac{1-(2c_{1})^{2}}{\varphi_{1}}-2nc_{1}\right]>0,$
$\mathrm{Ric}_{g_{\varphi_{1},\phi_{1}}}\left(X_{2},X_{2}\right)$ $=$ $\phi_{1}\left[-\phi_{1}^{\prime\prime}+(n-1)\frac{1-(\phi_{1}^{\prime})^{2}}{\phi_{1}}-2\frac{\phi_{1}^{\prime}\varphi_{1}^{\prime}}{\varphi_{1}}\right]$
$>$ $\phi_{1}\left(\frac{1}{2}-\frac{16c_{1}}{(9n)^{3}\sin(\frac{1}{10n})}\right)>0.$
For convenience, we can assume that there exists a constant $\xi\in(0,R_{1})$ such that
$\varphi_{1}(R_{1})=\phi_{1}(R_{1})=c_{1}(R_{1}+\xi).$
Part 2. For any $\alpha\leq\frac{c_{1}\epsilon}{10n}$, let $c_{2}=e^{-R_{1}^{2}\alpha^{-2}c_{1}^{-2}\epsilon^{-2}}$ and $\delta_{0}=c_{1}(R_{1}+\xi)(R_{1}+c_{2}^{-1})^{-\alpha}$.
Define $\varphi_{2}(r)=\delta_{0}(r+c_{2}^{-1})^{\alpha}$ on $[R_{1},\infty)$. Then we have
$\varphi^{\prime}_{2}(R_{1})=\alpha(R_{1}+c_{2}^{-1})^{-1}\varphi_{2}(R_{1})=\alpha c_{1}(R_{1}+\xi)(R_{1}+c_{2}^{-1})^{-1}\leq c_{1}=\varphi_{1}^{\prime}(R_{1}).$
Now we consider the Ricci curvature of $\mathrm{Ric}_{g_{\varphi_{2},\phi}}$. By a straightforward calculation, we have
$\mathrm{Ric}_{g_{\varphi_{2},\phi}}\left(X_{0},X_{0}\right)$ $=$ $-\left(2\frac{\varphi_{2}^{\prime\prime}}{\varphi_{2}}+n\frac{\phi^{\prime\prime}}{\phi}\right)$
$=$ $2\alpha(1-\alpha)(r+c_{2}^{-1})^{-2}-n\frac{\phi^{\prime\prime}}{\phi}\geq\alpha(r+c_{2}^{-1})^{-2}-n\frac{\phi^{\prime\prime}}{\phi},$
$\mathrm{Ric}_{g_{\varphi_{2},\phi}}\left(X_{1},X_{1}\right)$ $=$ $\varphi_{2}^{2}\left[-\frac{\varphi_{2}^{\prime\prime}}{\varphi_{2}}+\frac{1-(\varphi_{2}^{\prime})^{2}}{\varphi_{2}^{2}}-n\frac{\varphi_{2}^{\prime}\phi^{\prime}}{\varphi_{2}\phi}\right]$
$\geq$ $\varphi_{2}^{2}\left[\alpha(1-\alpha)(r+c_{2}^{-1})^{-2}+\frac{1}{2(r+c_{2}^{-1})^{2\alpha}}-\alpha n\frac{\phi^{\prime}}{(r+c_{2}^{-1})\phi}\right],$
$\mathrm{Ric}_{g_{\varphi_{2},\phi}}\left(X_{2},X_{2}\right)$ $=$ $\phi^{2}\left[-\frac{\phi^{\prime\prime}}{\phi}+(n-1)\frac{1-(\phi^{\prime})^{2}}{\phi^{2}}-2\frac{\varphi_{2}^{\prime}\phi^{\prime}}{\varphi_{2}\phi}\right]$
$=$ $\phi^{2}\left[-\frac{\phi^{\prime\prime}}{\phi}+(n-1)\frac{1-(\phi^{\prime})^{2}}{\phi^{2}}-2\alpha\frac{\phi^{\prime}}{(r+c_{2}^{-1})\phi}\right].$
Now we can find a positive function $\phi_{2}$ on $[R_{1},\infty)$ such that
[…]
Hence we have $1-(\phi^{\prime}_{2})^{2}\geq\frac{\epsilon}{2}$, $c_{1}\leq\phi^{\prime}_{2}\leq 1-\frac{\epsilon}{2}$, and $\phi\leq r$. It follows that when $r\geq R_{1}$, $\mathrm{Ric}_{g_{\varphi_{2},\phi_{2}}}>0$. Then we can find a constant $R>R_{1}$ such that $\phi_{2}(R)=(1-\epsilon)R+(1-\epsilon)c_{2}^{-1}$. Since $\phi^{\prime}_{2}>0$, we have $\phi^{\prime}_{2}(R)>1-\epsilon$.
Let $\varphi_{3}$, $\phi_{3}$ be non-negative functions on $[0,\infty)$ defined by
(20) $\varphi_{3}(r)=\left\{\begin{aligned} \varphi_{1}(r)\;\;\;\;&,\;\;\;\;r\leq R_{1},\\ \delta_{0}(r+c_{2}^{-1})^{\alpha}&,\;\;\;\;r\geq R_{1},\\ \end{aligned}\right.$
and
(21) $\phi_{3}(r)=\left\{\begin{aligned} \phi_{1}(r)\;\;\;\;&,\;\;\;\;r\leq R_{1},\\ \phi_{2}(r)\;\;\;\;&,\;\;\;\;R_{1}\leq r\leq R,\\ (1-\epsilon)(r+c_{2}^{-1})&,\;\;\;\;r\geq R.\\ \end{aligned}\right.$
By Lemma 3.2, there are functions $\tilde{\varphi}$ and $\tilde{\phi}$ such that $\mathrm{Ric}_{g_{\tilde{\varphi},\tilde{\phi}}}>0$, and for any $r\in\left[0,\frac{R_{1}}{2}\right]\cup[2R,\infty)$, we have $\tilde{\varphi}=\varphi_{3}$ and $\tilde{\phi}=\phi_{3}$. Consider the case $\phi=(1-\epsilon)r$ in (4)-(4), we see that there exists a constant $\tau_{1}=\tau_{1}(n,\epsilon,\alpha)>0$ such that $\mathrm{Ric}_{g_{\tilde{\varphi},\tilde{\phi}}}>\frac{\tau_{1}}{1+r^{2}}$, $\forall r\geq 2R$. Then $\mathrm{Ric}_{g_{\tilde{\varphi},\tilde{\phi}}}>0$ implies that there exists a constant $\tau_{2}=\tau_{2}(n,\epsilon,\alpha)>0$ such that $\mathrm{Ric}_{g_{\tilde{\varphi},\tilde{\phi}}}\geq\frac{\tau_{2}}{1+r^{2}}$, $\forall r\geq 0$.
Part 3. Now we consider the case $\delta<\delta_{0}$. Write $s=\delta_{0}^{-1}\delta\in(0,1)$. For any $0<t\leq c_{2}^{5}$, we define
(22) $\tilde{\varphi}_{t}(r)=\left\{\begin{aligned} t\sin(t^{-1}r)\;\;\;&,\;\;\;\;r\leq t\arccos(s),\\ s\tilde{\varphi}\left(r+\theta(t,s)\right)&,\;\;\;\;r\geq t\arccos(s),\\ \end{aligned}\right.$
where $\theta(t,s)=\arcsin\left(t\sqrt{s^{-2}-1}\right)-t\arccos(s)>0$.
It is easy to see that for any $t\geq 1$, $\mathrm{Ric}_{g_{\tilde{\varphi}_{t},\tilde{\phi}}}\geq\frac{1}{2t}$ for $r\leq t\arccos(s)$. When
$100\arccos(t\sqrt{s^{-2}-1})\leq r\leq\frac{1}{20n},$
one can apply Lemma 3.1 to show that
$\mathrm{Ric}_{g_{\tilde{\varphi}_{t},\tilde{\phi}}}\left(X_{0},X_{0}\right)$$\geq$ $2-n\frac{12r^{2}}{1+r^{4}}\geq 2-\frac{12}{10}\geq\frac{1}{2},$
$\mathrm{Ric}_{g_{\tilde{\varphi}_{t},\tilde{\phi}}}\left(X_{1},X_{1}\right)$$\geq$ $\tilde{\varphi}_{t}^{2}\left(1-n\frac{4r^{3}\cos\left(r+\theta(t,s)\right)}{\sin\left(r+\theta(t,s)\right)}\right)\geq\frac{\tilde{\varphi}_{t}^{2}}{2},$
$\mathrm{Ric}_{g_{\tilde{\varphi}_{t},\tilde{\phi}}}\left(X_{2},X_{2}\right)$$\geq$ $\tilde{\phi}^{2}\left[-12r^{2}+(n-1)\frac{1-16r^{6}}{2}-8r^{2}\right]$
$\geq$ $\tilde{\phi}^{2}\left(-\frac{12}{81}+\frac{1}{3}-\frac{8}{81}\right)\geq\frac{\tilde{\phi}^{2}}{81}.$
By Lemma 3.2, for any given $s,t$, we can modify $\tilde{\varphi}$ and $\tilde{\phi}$ near $t\arccos(k)$ such that they became smooth functions, and $\mathrm{Ric}_{g_{\tilde{\varphi}_{t},\tilde{\phi}}}\geq\frac{1}{100}$ for $r\leq\frac{1}{20n}$.
Choosing a cut-off function $\eta\in C^{\infty}([0,\infty))$ such that $0\leq\eta\leq 1$, $\eta(r)=1$ when $r\leq\frac{1}{100n}$, and $\eta(r)=0$ when $r\geq\frac{1}{50n}$. Now we can see that there exists a constant $t_{\delta}>0$ such that for any $t\in(0,2t_{\delta})$, we have $\mathrm{Ric}_{g_{\eta\tilde{\varphi}_{t}+(1-\eta)\tilde{\varphi},\tilde{\phi}}}\geq\frac{\tau_{2}}{2(1+r^{2})}$, $\forall r\geq 0$. Note that when $r\geq\frac{1}{200n}$, $\tilde{\varphi}_{t}$ converges to $\tilde{\varphi}$ as $t\to 0$ in the $C^{\infty}$-sense. Then we can prove this lemma by letting $\varphi=\eta\tilde{\varphi}_{t_{\delta}}+(1-\eta)\tilde{\varphi}$, $\phi=\tilde{\phi}$, $C=c_{2}^{-1}$, $R=c_{2}^{-10}$, and $\tau=c_{2}\tau_{2}$. ∎
As a corollary, we have:
Corollary 4.1.
Let $n\geq 2$. Then for any $0<\epsilon\leq\frac{1}{100}$, $0<\alpha\leq\alpha_{0}(n,\epsilon)$, $0<\kappa\leq\frac{1}{100}$, and $0<\delta\leq\delta_{0}(n,\epsilon,\alpha,\kappa)$, we can find constants $\tau=\tau(n,\epsilon,\alpha)\in(0,1)$, $0<\mu=\mu(\kappa|n,\epsilon,\alpha)=\Psi(\kappa|n,\epsilon,\alpha)$, a Riemannian manifold $(\mathcal{B},g)=(\mathcal{B}(\epsilon,\alpha,\delta,\kappa),g_{\epsilon,\alpha,\delta,\kappa})$, a compact totally geodesic submanifold $\mathcal{Z}\subset\mathcal{B}$ isometric to ${\mathbb{S}}_{\mu}^{n}$, and a diffeomorphism $F:B_{1}^{3}(0)\times{\mathbb{S}}^{n}\to\mathcal{B}$ such that:
  1. (1)
    $\mathrm{Ric}_{g}>\frac{\tau}{r+\kappa}$ on $\mathcal{B}$,
  2. (2)
    $F^{-1}(\mathcal{Z})=\{0\}\times{\mathbb{S}}^{n}$,
  3. (3)
    The pull-back $F^{*}g$ is given by $g_{\varphi,\phi}$ for some nonnegative functions satisfying (4) and (5),
  4. (4)
[…]
    $\left((\kappa,1)\times{\mathbb{S}}^{2}\times{\mathbb{S}}^{n},dr^{2}+\delta r^{\alpha}g_{{\mathbb{S}}^{2}}+(1-\epsilon)^{2}r^{2}g_{{\mathbb{S}}^{n}}\right).$
Moreover, if $(\mathcal{B},g)=(\mathcal{B}(\epsilon,\alpha,\delta,\kappa),g_{\epsilon,\alpha,\delta,\kappa})$ is the Riemannian manifold we constructed, then for any $t,\upsilon\in(0,1)$, there exists another metric $g_{t}$ on $\mathcal{B}$ such that $\mathrm{Ric}_{g_{t}}>\tau$ the pull-back $F^{*}g_{t}$ is given by $g_{\varphi_{t},\phi}$ for some nonnegative functions satisfying (4) and (5), and $\varphi_{t}=t\varphi$ on $[\upsilon,1)$.
Remark.
In this paper, $\Psi\left(\epsilon_{1},\cdots,\epsilon_{m}\big{|}u_{1},\cdots,u_{n}\right)$ denotes a function of $\epsilon_{1},\cdots,\epsilon_{m}$, $u_{1},\cdots,u_{n}$, such that for fixed $u_{i}$ we have $\lim_{\epsilon_{1},\cdots,\epsilon_{m}\to 0}\Psi=0$.
Proof.
By Lemma 4.1, for any positive $\epsilon\leq\frac{1}{100}$, $\alpha\leq\alpha_{0}(n,\epsilon)$ and $\delta_{1}\leq\delta_{0}(n,\epsilon,\alpha)$, we can find positive constants $R(n,\epsilon,\alpha)$, $C(n,\epsilon,\alpha)$, $\tau=\tau(n,\epsilon,\alpha)>0$ and functions $\varphi,\phi$ satisfying (4) and (5), such that $\mathrm{Ric}_{g_{\varphi_{1},\phi_{1}}}\geq\frac{\tau}{r^{2}+1}$, $\phi_{1}(0)=1$, $\phi_{1}|_{[R,\infty)}=(1-\epsilon)r$, $\varphi_{1}(0)=0$, and $\varphi_{1}|_{[R,\infty)}=\delta_{1}(r+C)^{\frac{\alpha}{2}}$.
Let $t\geq\frac{2R}{\kappa}$, $\varphi(r)=t^{-1}\varphi_{1}(tr)$ and $\phi(r)=t^{-1}\phi_{1}(tr)$. Then we have $\varphi(r)=t^{-1+\alpha}r^{\alpha}$ and $\phi(r)=(1-\epsilon)r$, $r\in[\kappa,\infty)$. Clearly, the rescaled metric $t^{-2}g_{\varphi_{1},\phi_{1}}$ can be represented by
$\left((0,1)\times{\mathbb{S}}^{2}\times{\mathbb{S}}^{n-1},dr^{2}+\varphi(r)^{2}g_{{\mathbb{S}}^{2}}+\phi(r)^{2}g_{{\mathbb{S}}^{n-1}}\right).$
It follows that
$\mathrm{Ric}_{g_{\varphi,\phi}}|_{r=r_{0}}=t^{2}\mathrm{Ric}_{g_{\varphi,\phi}}|_{r=tr_{0}}\geq\frac{t^{2}\tau}{t^{2}r_{0}^{2}+1}=\frac{\tau}{r_{0}^{2}+t^{-2}}\geq\frac{\tau}{r_{0}+\kappa},\;\;\forall r_{0}>0.$
Since $\varphi(0)=0$, $\varphi^{\prime}(0)=1$ and $\phi(0)=t^{-1}$, we see that $g_{\varphi,\phi}$ gives a Riemannian metric on $B_{1}^{3}(0)\times{\mathbb{S}}^{n}$, and the subset $\mathcal{Z}$ corresponding to $r=0$ is isometric to ${\mathbb{S}}_{t^{-1}}^{n}$. This completes the proof. ∎
By the last part of the proof of Lemma 4.1, one can obtain the following result.
Corollary 4.2.
Let $m=2$, $n\geq 2$, $\lambda,\mu>0$, and $\varphi,\phi$ be functions on $[0,\mu]$ satisfying (4), (5), and $\mathrm{Ric}_{g_{\varphi,\phi}}>\lambda$. Assume that there exists an open neighborhood $U$ of $0\in[0,r]$ such that $\varphi(r)=A^{-1}\sin(Ar)$ and $\phi(r)=A^{-1}+A^{3}r^{4}$ on $U$. Then for any $\delta,\epsilon\in(0,1)$, we can find a function $\tilde{\phi}\geq 0$ satisfying (4) , $\mathrm{Ric}_{g_{\varphi,\tilde{\phi}}}>\lambda$, $|\phi-\tilde{\phi}|\leq\epsilon$, and $\tilde{\phi}=\delta\phi$ on $\left[\frac{\mu}{2},\mu\right]$.
5. Construction of $(M_{i},g_{i})$
In this section, we will construct the sequence of $(n+2)$-dimensional manifolds $(M_{i},g_{i})\stackrel{{\scriptstyle GH}}{{\longrightarrow}}(X_{\epsilon},d_{\epsilon})$ for some given $\epsilon$. Our construction is based on an induction process.
5.1. Puncture lemma
At first, in order to make the first hole on $X$, we need following lemma. This lemma is a combination of Lemma 4.1 and Lemma 3.3. Since we plan to use this lemma to the interior of a manifold with boundary, we need to consider non-compact manifolds here.
Lemma 5.1.
Let $(M_{B},g_{B})$ be an $n$-dimensional Riemannian manifold (need not be complete), $n\geq 3$, $p\in M_{B}$, $\mathcal{U}_{p}$ be a pre-compact open neighborhood of $p$, and $f$ be a positive smooth function on $\bar{\mathcal{U}}_{p}$. Assume that the warped product space $(\mathcal{U}_{p}\times{\mathbb{S}}^{2},g)$ with metric $g=g_{B}+f^{2}g_{{\mathbb{S}}^{2}}$ satisfying that $\mathrm{Ric}_{g}>\lambda\in\mathbb{R}$. Then there exist constants $r_{0}\in(0,1]$, $\epsilon_{0}\in(0,\frac{1}{100})$ depends on $(\mathcal{U}_{p}\times{\mathbb{S}}^{2},g)$, such that for any $0<\epsilon<\epsilon_{0}$, $0<r^{*}\leq r^{*}_{0}(g,\epsilon)$, and $0<\delta\leq\delta_{0}(g,\epsilon,r^{*})$, there exists an $(n+2)$-dimensional Riemannian manifold $(\mathcal{V}_{p},\hat{{g}})$, satisfying the following properties:
  1. (i).
    The closed ball $\bar{B}^{g_{B}}_{4r_{0}}(p)\subset(M_{B},g_{B})$ is a compact subset of $\mathcal{U}_{p}$,
  2. (ii).
    $\mathrm{Ric}_{\hat{g}}>\lambda$ on $\mathcal{V}_{p}$.
Moreover, there exist an $n$-dimensional compact totally geodesic submanifold $\mathcal{Z}\subset\mathcal{V}_{p}$ isometric to ${\mathbb{S}}_{r^{*}}^{n}$, and a diffeomorphism $F:\mathcal{V}_{p}\backslash\mathcal{Z}\to(\mathcal{U}_{p}\backslash\bar{B}^{g_{B}}_{r^{*}}(p))\times{\mathbb{S}}^{2}$, such that:
  1. (iii).
    The pull-back of metric $(F^{-1})^{*}\hat{g}$ is given by the warped product $\left((\mathcal{U}_{p}\backslash\bar{B}^{g_{B}}_{r^{*}}(p))\times{\mathbb{S}}^{2},\hat{{g}}_{B}+\hat{f}^{2}\right)$,
  2. (iv).
    $\|\hat{f}\|_{L^{\infty}(B^{g_{B}}_{3r_{0}}(p))}\leq\delta\|f\|_{L^{\infty}(B^{g_{B}}_{3r_{0}}(p))}$,
  3. (v).
    $\hat{{g}}_{B}=g_{B}$ and $\hat{f}=\delta^{2}f$ on $\mathcal{U}_{p}\backslash B^{g_{B}}_{2r_{0}}(p)$.
  4. (vi).
    The Gromov-Hausdorff distance $d_{GH}\left((\mathcal{U}_{p}\backslash\bar{B}^{g_{B}}_{r^{*}}(p),d_{\hat{{g}}_{B}}),(\mathcal{U}_{p},d_{g_{B}})\right)\leq\epsilon$,
  5. (vii).
    The complement of $({B}^{g_{B}}_{2r^{*}}(p)\backslash\bar{B}^{g_{B}}_{r^{*}}(p),\hat{{g}}_{B})$ is a smooth Riemannian manifold with boundary,
  6. (viii).
    The continuous extension of $\pi_{\mathcal{U}_{p}}\circ F:\pi_{\mathcal{U}_{p}}\circ F:(\mathcal{V}_{p}\backslash\mathcal{Z},\hat{g})\to\left(\mathcal{U}_{p}\backslash\bar{B}^{g_{B}}_{r^{*}}(p),\hat{{g}}_{B}\right)$ gives an isometric $\mathcal{Z}\to\partial{B}^{\hat{{g}}_{B}}_{r^{*}}(p)$,
  7. (ix).
    The restriction of inclusion map $\text{Id}:\left(\mathcal{U}_{p}\backslash\bar{B}_{r^{*}}(p),d_{\hat{{g}}_{B}}\right)\to\left(\mathcal{U}_{p},d_{g_{B}}\right)$ on $\mathcal{U}_{p}\backslash\bar{B}^{g_{B}}_{r_{0}}(p)$ is $(1+\epsilon)$-bi-Lipschitz.
Remark.
As we pointed out in Corollary 4.1, if $(\mathcal{V}_{p},\hat{g})$ is the Riemannian manifold we constructed in this lemma, then for any $t\in(0,1)$ and any open neighborhood $\mathcal{U}_{\mathcal{Z}}$ of $\mathcal{Z}$, there exists another metric $\hat{g}_{t}$ on $\mathcal{V}_{p}$ such that $\mathrm{Ric}_{\hat{g}_{t}}>\tau$, and the pull-back $(F^{-1})^{*}\hat{g}_{t}$ is given by the warped product $\hat{g}_{B}+\hat{f}_{t}^{2}$, and $\hat{f}_{t}=t\hat{f}$ outside $F^{-1}(\mathcal{U}_{\mathcal{Z}})$.
Proof.
Our construction is divided into three parts. The first step is to add a conical singularity to the manifold $M_{B}$ using Lemma 3.3. The second step is to change the conical singularity we added to a hole using Lemma 4.1 or Corollary 4.1. In the first two parts, we will prove (i) and (ii). In the last part, we will construct the map $F$ and prove the properties (iii)-(ix).
Part 1. By scaling, one can find a small constant $r_{B}=r_{B}(g_{B})>0$ such that $\bar{B}^{g}_{10r_{B}}(p)$ is a compact subset of $\mathcal{U}_{p}$, and the manifold $\left(B^{r^{-2}_{B}g_{B}}_{2}(p)\times{\mathbb{S}}^{2},r^{-2}_{B}g\right)=\left(B^{g}_{2r_{B}}(p)\times{\mathbb{S}}^{2},r_{B}^{-2}g_{B}+r_{B}^{-2}f^{2}g_{{\mathbb{S}}^{2}}\right)$ satisfying the conditions (11)-(13). Note that for any $\lambda\in(0,\infty)$, the data $(g_{B},f)$ satisfies (13) if and only if $(g_{B},\lambda f)$ satisfies (13). By replaying $g$ by $r^{-2}_{B}g$, we can assume that $\bar{B}^{g_{B}}_{10}(p)$ is a compact subset of $\mathcal{U}_{p}$, and the manifold $\left(B^{g_{B}}_{2}(p)\times{\mathbb{S}}^{2},g\right)$ satisfying the conditions in Lemma 3.3. Then one can obtain (i) by choosing $r_{0}=1$. Without loss of generality, we can assume that $\mathrm{Ric}_{g}\geq\lambda+\mu$ for some constant $\mu>0$ on $\bar{B}^{g_{B}}_{10}(p)\times{\mathbb{S}}^{2}$.
[…]
Now we will describe how to select the data $(\epsilon_{1},\alpha_{1},\hat{r}_{1},\hat{\delta}_{1})$. By Lemma 3.3, there exists a constant $C=C(n)>0$ such that the Ricci lower bound $\mathrm{Ric}_{\hat{g}_{1}}>(\lambda+\mu)-C(n)\epsilon_{1}$ holds for $\frac{\hat{r}_{1}}{2}\leq r\leq 2$. Hence we can conclude that $\mathrm{Ric}_{\hat{g}_{1}}\geq\lambda+\frac{\mu}{2}$ holds for $\frac{\hat{r}_{1}}{2}\leq r\leq 2$ if $\epsilon_{1}\leq\frac{\mu}{10C}$. Then we choose $\epsilon_{1}\leq\frac{\mu}{20C}$, and $\alpha_{1}\leq\alpha(n,|\lambda|)$ such that the identity map $\text{Id}:(B_{2}(p),\hat{g}_{B,1})\to(B_{2}(p),g_{B})$ is $(1+\epsilon^{2})$-bi-Lipschitz.
Now we can apply Lemma 3.3 again to show that $\hat{g}_{1}=dr^{2}+(1-\epsilon_{1})^{2}r^{2}g_{{\mathbb{S}}^{n}}+\delta_{1}^{2}r^{2\alpha}g_{{\mathbb{S}}^{2}}$, $\forall r\leq\hat{r}_{1}$, where $\delta_{1}=\Psi\left(\hat{\delta}_{1}\|f_{0}\|_{L^{\infty}}\bigg{|}n,\alpha_{1},\lambda,\epsilon_{1}\right)$. Now we move on to the next step, which is to make a hole at $p$. We will determine the constants $\hat{r}_{1}$ and $\hat{\delta}_{1}$ in the last part.
Part 2. By Corollary 4.1, for any constants $\kappa_{1}\leq\kappa_{0}(n,\epsilon_{1},\alpha_{1})<1$ and $\delta^{\prime}_{1}\leq\delta^{\prime}_{0}(n,\epsilon_{1},\alpha_{1},\kappa_{1})<1$, then there exist a constant $\tau(n,\epsilon_{1},\alpha_{1})>0$, a Riemannian manifold $(\mathcal{B},g_{\mathcal{B}})$, a compact submanifold $\mathcal{Z}\subset\mathcal{B}$ isometric to ${\mathbb{S}}_{\sigma}^{n}$, and a diffeomorphism $F_{\mathcal{B}}:B_{1}^{3}(0)\times{\mathbb{S}}^{n}\to\mathcal{B}$ such that:
  • $0<\sigma=\Psi(\kappa_{1}|n,\epsilon_{1},\alpha_{1})$,
  • $\mathrm{Ric}_{g_{\mathcal{B}}}>\frac{\tau}{r+\kappa_{1}}$ on $\mathcal{B}$,
  • $F_{\mathcal{B}}^{-1}(\mathcal{Z})=\{0\}\times{\mathbb{S}}^{n}$,
  • The pull-back $F_{\mathcal{B}}^{*}g_{\mathcal{B}}$ is given by $g_{\varphi,\phi}$ for some nonnegative functions satisfying (4) and (5),
[…]
    $\left((\kappa_{1},1)\times{\mathbb{S}}^{2}\times{\mathbb{S}}^{n-1},dr^{2}+\delta^{\prime}_{1}r^{\alpha_{1}}g_{{\mathbb{S}}^{2}}+(1-\epsilon_{1})^{2}r^{2}g_{{\mathbb{S}}^{n-1}}\right).$
Let $\hat{r}_{1}=10\kappa_{1}$ and $\delta^{\prime}_{1}=\delta_{1}$. Now we can glue $(F_{\mathcal{B}}(B_{2\kappa}^{3}(0)\times{\mathbb{S}}^{n}),g_{\mathcal{B}})$ to $\left(B_{2}(p)\times{\mathbb{S}}^{2},\hat{g}\right)$ by the identity map on $(4\kappa_{1},5\kappa_{1})\times{\mathbb{S}}^{2}\times{\mathbb{S}}^{n-1}$.
Let $(\mathcal{V}_{p},\hat{{g}})$ denote the manifold of $(F_{\mathcal{B}}(B_{2\kappa}^{3}(0)\times{\mathbb{S}}^{n}),g_{\mathcal{B}})$ glued to $\left(B_{2}(p)\times{\mathbb{S}}^{2},\hat{g}\right)$. Since $\tau$ is independent of $\kappa_{1}$ and $\delta^{\prime}_{1}$, one can see there exists a constant $\kappa_{2}=\kappa_{0}(n,\epsilon_{1},\alpha_{1},|\lambda|)$ such that when $\kappa_{1}\in(0,\kappa_{2})$, $\mathrm{Ric}_{g_{\mathcal{B}}}>\frac{\tau}{r+\kappa_{1}}>\lambda$, $\forall r\leq\hat{r}_{1}$. Then we can conclude that the manifold $(\mathcal{V}_{p},\hat{{g}})$ satisfies (ii) when $\hat{r}_{1}\in(0,\frac{\kappa_{2}}{10})$.
Part 3. Let $(B_{\mathcal{B}},g_{B_{\mathcal{B}}})=((0,1)\times{\mathbb{S}^{n}},dr^{2}+\phi^{2}g_{{\mathbb{S}}^{n-1}})$. Then we have $\phi=(1-\epsilon_{1})r$ for $r\geq\kappa_{1}$. Then we can glue $(B_{\mathcal{B}},g_{B_{\mathcal{B}}})$ to $\left(B_{2}(p),\hat{g}_{B,1}\right)$ by the identity map on $(4\kappa_{1},5\kappa_{1})\times{\mathbb{S}}^{n-1}$. Let $(B_{\mathcal{V}},\hat{{g}}_{B_{\mathcal{V}}})$ denote the manifold of $(B_{\mathcal{B}},g_{B_{\mathcal{B}}})$ glued to $\left(B_{2}(p),\hat{g}_{B,1}\right)$. It is easy to see that there is a natural projection $\pi_{\mathcal{V}}:\mathcal{V}\backslash\mathcal{Z}\to B_{\mathcal{V}}$, and this projection gives a warped product structure $(B_{\mathcal{V}}\times{\mathbb{S}}^{2},\hat{{g}}_{B_{\mathcal{V}}}+\hat{f}_{B_{\mathcal{V}}}^{2}g_{{\mathbb{S}}^{2}})$ corresponding to the ${\mathbb{S}}^{2}$-bundle $\pi_{\mathcal{V}}:\mathcal{V}\backslash\mathcal{Z}\to B_{\mathcal{V}}$.
Now we construct a Lipschitz map
$\varrho:B_{\mathcal{B}}=\left((0,1)\times{\mathbb{S}}^{n-1},dr^{2}+\phi^{2}g_{{\mathbb{S}}^{n-1}}\right)\to\left((0,1)\times{\mathbb{S}}^{n-1},dr^{2}+(1-\epsilon_{1})^{2}r^{2}g_{{\mathbb{S}}^{n-1}}\right)\subset B_{2}(p)$
by letting $\varrho(r,x)=((1-\epsilon_{1})^{-1}\phi(r),x)$. It is easy to see that for any given $r^{\prime}\in(0,1)$, the restriction $\varrho|_{r=r^{\prime}}$ is isometric. Let $r^{*}=(1-\epsilon)^{-1}\phi(0)=\Psi(\kappa|\epsilon_{1},\alpha)$. Then we can get a diffeomorphism $B_{\mathcal{V}}\to\mathcal{U}_{p}\backslash\bar{B}^{g_{B}}_{r^{*}}(p)$ by gluing $\varrho$ and the identity map of $\mathcal{U}_{p}\backslash\bar{B}^{g_{B}}_{3\kappa_{1}}(p)$. Hence one can construct the map $F$ by combining this diffeomorphism and the warped product structure we constructed above. By the construction of $F$, one can obtain (iii)-(viii).
It is sufficient to prove (ix) now.
Without loss of generality, we assume that $r^{*}\leq\frac{r_{0}}{10}$. Let $x,y\in\mathcal{U}_{p}\backslash\bar{B}^{g_{B}}_{r_{0}}(p)$. When $d_{\hat{g}_{B,1}}(x,y)\leq r_{0}-10\kappa_{1}$, the intersection of $\bar{B}^{g_{B}}_{\kappa_{1}}(p)$ and the geodesic between $x,y$ in $(\mathcal{U}_{p},\hat{g}_{B,1})$ is empty, and $d_{\hat{g}_{B,1}}(x,y)=d_{\hat{g}_{B}}(x,y)$. If $d_{\hat{g}_{B,1}}(x,y)\geq r_{0}-10\kappa_{1}$, then we have
$d_{\hat{g}_{B}}(x,y)-10\kappa_{1}\leq d_{\hat{g}_{B,1}}(x,y)\leq d_{\hat{g}_{B}}(x,y)+10\kappa_{1}.$
It follows that the restriction of inclusion map $\text{Id}:\left(\mathcal{U}_{p}\backslash\bar{B}_{r^{*}}(p),d_{\hat{{g}}_{B}}\right)\to\left(\mathcal{U}_{p},d_{\hat{g}_{B,1}}\right)$ on $\mathcal{U}_{p}\backslash\bar{B}^{g_{B}}_{r_{0}}(p)$ is $(1+\Psi(\kappa_{1}|n,r_{0}))$-bi-Lipschitz. Since the identity map $\text{Id}:(B_{2}(p),\hat{g}_{B,1})\to(B_{2}(p),g_{B})$ is $(1+\epsilon^{2})$-bi-Lipschitz, we see that (ix) holds when $\kappa_{1}$ is sufficiently small.
Now we describe the choice of $\hat{r}_{1}$ and $\hat{\delta}_{1}$ in the first part. Since $\hat{r}_{1}=10\kappa_{1}$, we can determine $\hat{r}_{1}$ by choosing surfficiently small $\kappa_{1}>0$. Note that we can always shrink $\hat{\delta}_{1}$ arbitrarily, so we just need to take sufficiently small $\hat{\delta}_{1}$ such that $\delta^{\prime}_{1}=\delta_{1}$ is sufficiently small, and all the above arguments hold. This completes the proof. ∎
5.2. Puncturing the first hole
Now we begin to construct the sequence of $(n+2)$-dimensional manifolds $(M_{i},g_{i})\stackrel{{\scriptstyle GH}}{{\longrightarrow}}(X_{\epsilon},d_{\epsilon})$.
Let $(M_{0},g_{0})$ denotes the complete Riemannian manifold $\left(X\times{\mathbb{S}}^{2},h+\frac{1}{1+|\lambda|^{2}}g_{{\mathbb{S}}^{2}}\right)$. Since $\mathrm{Ric}_{h}>\lambda$, one can see that $\mathrm{Ric}_{g_{0}}>\lambda$.
Let $p_{0}\in X$, $p_{1}\in B^{h}_{\frac{7}{4}}(p_{0})\backslash\bar{B}^{h}_{\frac{5}{4}}(p_{0})$, and $f_{0}=\frac{r_{1}^{-2}}{1+|\lambda|^{2}}$. Fix $r_{1}\leq\frac{\epsilon}{100}$, $\epsilon_{1}\leq\frac{\epsilon}{100}$, $r^{*}_{1}\leq\frac{r_{1}}{100}$ and $\delta_{1}\leq\frac{\epsilon_{1}}{100}$ satisfying the conditions in Lemma 5.1. Then one can apply Lemma 5.1 to the data $(p_{1},\epsilon_{1},r^{*}_{1},\delta_{1})$ and construct:
  • An $(n+2)$-dimensional Riemannian manifold $(M_{1},g_{1})$ with $\mathrm{Ric}_{g_{1}}>\lambda$,
  • An $n$-dimensional compact totally geodesic submanifold $\mathcal{Z}_{1}\subset M_{1}$ isometric to ${\mathbb{S}}_{r_{1}^{*}}^{n}$,
  • A Riemannian metric $g_{X,1}$ and a positive function $f_{1}$ on $X\backslash\bar{B}^{h}_{r_{1}^{*}}(p_{1})$,
  • An isomorphism $F_{1}:(M_{1}\backslash\mathcal{Z}_{1},g_{1})\to\left((X\backslash\bar{B}^{h}_{r_{1}^{*}}(p_{1}))\times{\mathbb{S}}^{2},g_{X,1}+f_{1}^{2}g_{{\mathbb{S}}^{2}}\right)$,
such that:
  • On $X\backslash B^{h}_{r_{1}}(p_{1})$, $f_{1}=\delta_{1}f_{0}$ and $g_{X,1}=h$,
  • On $B^{h}_{2r_{1}}(p_{1})$, $|f_{1}|\leq\epsilon_{1}$,
  • The continuous extension of $F_{1}$ gives an isometric $Z\to\partial B^{h}_{r^{*}_{1}}(p_{1})$,
  • The Gromov-Hausdorff distance $d_{GH}\left((X\backslash\bar{B}^{h}_{r_{1}^{*}}(p_{1}),g_{X,1}),(X,h)\right)\leq\epsilon_{1}$.
Let $(X_{1},d_{1})$ be the completion of $(X\backslash\bar{B}^{h}_{r_{1}^{*}}(p),g_{X,1})$. Then we see that the metric space $(X_{1},d_{1})$ is given by a smooth Riemannian manifold with boundary $\partial X_{1}\cong{\mathbb{S}}_{r_{1}^{*}}^{n}$.
5.3. Induction process
Our next step is to construct $(M_{i},g_{i})$ by induction. We state here the inductive hypothesis we need for induction on $i$.
Inductive Hypothesis (about $i$).
Fix $p_{0}\in X$. Now we assume that we have constructed:
  1. 1)i.
    An $(n+2)$-dimensional Riemannian manifold $(M_{i},g_{i})$ with $\mathrm{Ric}_{g_{i}}>\lambda$,
  2. 2)i.
    A pre-compact open subset $\mathcal{W}_{i}=\bigsqcup\limits_{j=1}^{N_{i}}\mathcal{U}_{j}\subset\bigsqcup\limits_{j=-i}^{i}B^{h}_{2^{j+1}-2^{-3i}}(p_{0})\backslash\bar{B}^{h}_{2^{j}+2^{-3i}}(p_{0})\subset(X,h)$,
  3. 3)i.
    An $n$-dimensional compact totally geodesic submanifold $\mathcal{Y}_{i}=\bigsqcup\limits_{j=1}^{N_{i}}\mathcal{Z}_{j}\subset M_{i}$,
  4. 4)i.
    A Riemannian metric $g_{X,i}$ and a positive function $f_{i}$ on $X\backslash\mathcal{W}_{i}$,
  5. 5)i.
    An isomorphism $F_{i}:(M_{i}\backslash\mathcal{Y}_{i},g_{i})\to\left((X\backslash\bar{\mathcal{W}}_{i})\times{\mathbb{S}}^{2},g_{X,i}+f_{i}^{2}g_{{\mathbb{S}}^{2}}\right)$,
such that:
  1. 6)i.
    The function $|f_{i}|\leq 2^{-i}$,
  2. 7)i.
    The closures $\bar{\mathcal{U}}_{j}$ are diffeomorphic to the closed unit $n$-ball,
  3. 8)i.
    The completion of $\left(X\backslash\bar{\mathcal{W}}_{i},d_{g_{X,i}}\right)$ is a smooth Riemannian manifold with boundary $\partial\mathcal{W}_{i}$,
  4. 9)i.
    The metric $g_{X,i}=h$ on $B^{h}_{2^{-i}}(p_{0})\cup\left(X\backslash B^{h}_{2^{i+1}}(p_{0})\right)\cup\bigsqcup\limits_{j=-i}^{i}B^{h}_{2^{j}+2^{-4i}}(p_{0})\backslash\bar{B}^{h}_{2^{j}-2^{-4i}}(p_{0})\subset(X,h)$,
  5. 10)i.
    The Gromov-Hausdorff distance $d_{GH}\left((X\backslash\bar{\mathcal{W}}_{i}),d_{g_{X,i}}),(X,h)\right)\leq\left(1-2^{-i}\right)\epsilon$,
  6. 11)i.
    For any $x\in B^{h}_{2^{i}}(p_{0})\backslash\bar{\mathcal{W}}_{i}\subset\left(X\backslash\bar{\mathcal{W}}_{i},d_{g_{X,i}}\right)$, the distance $d_{g_{X,i}}(x,\mathcal{W}_{i})\leq 2^{2-i}$,
  7. 12)i.
    For any $x\in(B^{h}_{2^{i}}\backslash\bar{\mathcal{W}}_{i},d_{g_{X,i}})$, there exist a constant $c_{i}\in(0,2^{-i})$ and a closed subset $K_{x,i}$ of $\mathbb{R}^{n}$ satisfies that $c_{i}^{-1}d_{GH}\left(B^{g_{X,i}}_{c_{i}}(x),K_{x,i}\right)\leq 2^{1-i}(1-2^{-i})$.
Note that the last of these property listed in Proposition 5.1 derives from the following fact.
Fact 5.1.
Let $(M,g)$ be a smooth Riemannian manifold with boundary, and $\Omega$ is a compact subset of $M$. Then for any $x\in\Omega$ and $\epsilon>0$, there exists a constant $r_{0}=r_{0}(\Omega,g)>0$ such that for any $0<r<r_{0}$, we can find a subset $K_{x,r}$ of $\mathbb{R}^{n}$ satisfies that
$r^{-1}d_{GH}\left(B^{g}_{r}(x),K_{x,r}\right)\leq\epsilon.$
Now we construct the manifold $(M_{i+1},g_{i+1})$.
Let us begin with choosing a discrete subset $P=\{p_{k}\}_{k=1}^{N^{\prime}_{i}}$ of $X\backslash\bar{\mathcal{W}}_{i}$ such that for any $x\in B^{h}_{2^{i}}(p_{0})\backslash\mathcal{W}_{i}\subset\left(X\backslash\bar{\mathcal{W}}_{i},d_{g_{X,i}}\right)$, the distance $d_{g_{X,i}}(x,\mathcal{W}_{i}\cup P)+d_{h}(x,\mathcal{W}_{i}\cup P)\leq 2^{-2-i}$.
Then we can construct $(M_{i+1},g_{i+1})$, $\mathcal{W}_{i+1}$, $\mathcal{Y}_{i+1}$, $\left((X\backslash\mathcal{W}_{i+1})\times{\mathbb{S}}^{2},g_{X,i+1}+f_{i+1}^{2}g_{{\mathbb{S}}^{2}}\right)$ and $F_{i+1}$ by applying Lemma 5.1 to each point in the discrete set $P=\{p_{k}\}_{k=1}^{N^{\prime}_{i}}\subset\left(X\backslash\bar{\mathcal{W}}_{i},d_{g_{X,i}}\right)$. We will finish it by induction on $k$. We state here the inductive hypothesis we need for induction on $i$.
Inductive Hypothesis (about $k$).
Fix $p_{0}\in X$. Now we assume that we have constructed:
  1. 1)i,k.
    An $(n+2)$-dimensional Riemannian manifold $(M_{i,k},g_{i,k})$ with $\mathrm{Ric}_{g_{i,k}}>\lambda$,
  2. 2)i,k.
    A pre-compact open subset
    $\mathcal{W}_{i,k}=\left(\bigsqcup\limits_{j=1}^{N_{i}}\mathcal{U}_{j}\right)\bigsqcup\left(\bigsqcup\limits_{l=1}^{k}\mathcal{U}^{\prime}_{l}\right)\subset\bigsqcup\limits_{j=-i-1}^{i+1}B^{h}_{2^{j+1}-2^{-3i-3}}(p_{0})\backslash\bar{B}^{h}_{2^{j}+2^{-3i-3}}(p_{0})\subset(X,h),$
  3. 3)i,k.
    An $n$-dimensional compact totally geodesic submanifold $\mathcal{Y}_{i,k}=\left(\bigsqcup\limits_{j=1}^{N_{i}}\mathcal{Z}_{j}\right)\bigsqcup\left(\bigsqcup\limits_{l=1}^{k}\mathcal{Z}^{\prime}_{l}\right)\subset M_{i,k}$,
  4. 4)i,k.
    A Riemannian metric $g_{X,i,k}$ and a positive function $f_{i,k}$ on $X\backslash\mathcal{W}_{i,k}$,
  5. 5)i,k.
    An isomorphism $F_{i,k}:(M_{i,k}\backslash\mathcal{Y}_{i,k},g_{i,k})\to\left((X\backslash\bar{\mathcal{W}}_{i,k})\times{\mathbb{S}}^{2},g_{X,{i,k}}+f_{i,k}^{2}g_{{\mathbb{S}}^{2}}\right)$,
such that:
  1. 6)i,k.
    The closures $\bar{\mathcal{U}}_{j}$ and $\bar{\mathcal{U}}^{\prime}_{l}$ are diffeomorphic to the closed unit $n$-ball,
  2. 7)i,k.
    The completion of $\left(X\backslash\bar{\mathcal{W}}_{i,k},d_{g_{X,i,k}}\right)$ is a smooth Riemannian manifold with boundary $\partial\mathcal{W}_{i,k}$,
  3. 8)i,k.
    The metric $g_{X,i,k}=h$ on
    $B^{h}_{2^{-i-1}(1+2^{-3i-3-k})}(p_{0})\cup\left(X\backslash B^{h}_{2^{i+2}(1+2^{-5i-3-k})}(p_{0})\right)$
[…]
  4. 9)i,k.
    For any $x\in(B^{h}_{2^{i}}\backslash\bar{\mathcal{W}}_{i},d_{g_{X,i}})$, there exist a constant $c_{i}\in(0,2^{-i})$ and a closed subset $K_{x,i}$ of $\mathbb{R}^{n}$ satisfies that $c_{i}^{-1}d_{GH}\left(B^{g_{X,i}}_{c_{i}}(x),K_{x,i}\right)\leq 2^{-i}(2-2^{-i}+k2^{-i-{N^{\prime}_{i}}})$,
  5. 10)i,k.
    For any point $p_{l}\in P$, $p_{l}\in\bar{\mathcal{U}}^{\prime}_{l^{\prime}}$ if and only if $l=l^{\prime}$.
Now we can apply Lemma 5.1 around $p_{k+1}$ to get a smooth Riemannian manifold $(\mathcal{V}_{k+1},g_{\mathcal{V}})$, an open subset $\mathcal{V}^{\prime}_{k+1}$ of $\mathcal{V}_{k+1}$, preconpact open neighborhoods $\mathcal{U}^{\prime\prime\prime}_{k+1}\subset\subset\mathcal{U}^{\prime\prime}_{k+1}$ of $p_{k+1}$ in $X\backslash\bar{\mathcal{W}}_{i,k}$, such that $\mathcal{V}_{k+1}\backslash\mathcal{V}^{\prime}_{k+1}$ is compact, and $(\mathcal{V}^{\prime}_{k+1},g_{\mathcal{V}})$ is isometric to $((\mathcal{U}^{\prime\prime}_{k+1}\backslash\mathcal{U}^{\prime\prime\prime}_{k+1})\times{\mathbb{S}}^{2},g_{X,i,k}+\delta^{2}f_{i,k}^{2}g_{{\mathbb{S}}^{2}})$, where $\delta\in(0,1)$ is a constant.
By Corollary 4.2, one can obtain $f_{i,k+1}$ by modifying $\delta f_{i,k}$ on a small open neighborhood of $\mathcal{Y}_{i,k}$ such that the completion of $g_{X,i,k}+f_{i,k+1}^{2}g_{{\mathbb{S}}^{2}}$ gives a smooth Riemannian metric around $\mathcal{Y}_{i,k}$. Then we can construct $(M_{i,k+1},g_{i,k+1})$ by gluing $(\mathcal{V}_{k+1},g_{\mathcal{V}})$ to $\left(M_{i,k}\backslash\left(\mathcal{U}^{\prime\prime\prime}_{k+1}\times{\mathbb{S}}^{2}\right),g_{X,i,k}+\delta^{2}f_{i,k}^{2}g_{{\mathbb{S}}^{2}}\right)$. By applying Lemma 5.1 again, one can obtain $\mathcal{U}^{\prime}_{k+1}$, $\mathcal{Z}^{\prime}_{k+1}$, $g_{X,i,k+1}$, $f_{i,k+1}$ and $F_{i,k+1}$ satisfying the conditions 1)i,k-10)i,k.
Note here that we can ensure that for any $p\in\{p_{0}\}\bigsqcup\left(\bigsqcup\limits_{i=0}^{\infty}\partial B^{h}_{2^{-i}}(p_{0})\right)\bigsqcup\partial\mathcal{W}_{i,k+1}$ and $\sigma\in(0,1)$, the Gromov-Hausdorff distance
$\sigma^{-1}d_{GH}(B_{\sigma}^{g_{X,i,k}}(p),B_{\sigma}^{g_{X,i,k+1}}(p+1))\leq 2^{-101(i+k)}$
by choosing sufficiently small $r^{*},\delta>0$ when we applying Lemma 5.1. Similarly, by choosing sufficiently small $r^{*},\delta>0$, one can obtain that the identity map of $B^{h}_{2^{-2-i}}(p_{0})$ is also $(1+\Psi(r^{*},\delta|g_{X,i}))$-bi-Lipschitz.
Now we finish the induction process about $k$.
Let $(M_{i+1},g_{i+1},g_{X,i+1},f_{i+1})=(M_{i,N^{\prime}_{i}},g_{i,N^{\prime}_{i}},g_{X,i,N^{\prime}_{i}},f_{i,N^{\prime}_{i}})$, $\mathcal{W}_{i+1}=\mathcal{W}_{i,N^{\prime}_{i}}$, $\mathcal{Y}_{i+1}=\mathcal{Y}_{i,N^{\prime}_{i}}$ and $F_{i+1}=F_{i,N^{\prime}_{i}}$. Thus we can obtain the following property.
Proposition 5.1.
Let $(X,h)$ be an $n$-dimensional smooth complete Riemannian manifold with $n\geq 3$ and $\mathrm{Ric}_{h}>\lambda\in\mathbb{R}$. Fix $p_{0}\in X$ and $\epsilon>0$. Then we have:
  • A sequence of numbers $c_{i}\to\infty$ as $i\to\infty$,
  • A sequence of $(n+2)$-dimensional Riemannian manifold $(M_{i},g_{i})$ with $\mathrm{Ric}_{g_{i}}>\lambda$,
  • An increasing sequence of open subsets
    $\mathcal{W}_{i}=\bigsqcup\limits_{j=1}^{N_{i}}\mathcal{U}_{j}\subset\bigsqcup\limits_{j=-i}^{i}B^{h}_{2^{j+1}-2^{-3i}}(p_{0})\backslash\bar{B}^{h}_{2^{j}+2^{-3i}}(p_{0})\subset(X,h),$
  • A sequence of $n$-dimensional compact submanifolds $\mathcal{Y}_{i}=\bigsqcup_{j=1}^{N_{i}}\mathcal{Z}_{j}\subset M_{i}$,
  • A sequence of Riemannian metrics $g_{X,i}$ and a sequence of positive functions $f_{i}$ on $X\backslash\mathcal{W}_{i}$,
  • A sequence of isomorphisms $F_{i}:(M_{i}\backslash\mathcal{Y}_{i},g_{k})\to\left((X\backslash\bar{\mathcal{W}}_{i})\times{\mathbb{S}}^{2},g_{X,i}+f_{i}^{2}g_{{\mathbb{S}}^{2}}\right)$,
such that:
  • The functions $|f_{i}|\leq 2^{-i}$,
  • The closures $\bar{\mathcal{U}}_{j}$ are diffeomorphic to the closed unit $n$-ball,
  • The completions of $\left(X\backslash\bar{\mathcal{W}}_{i},d_{g_{X,i}}\right)$ are smooth Riemannian manifolds with boundary $\partial{\mathcal{W}_{i}}$,
  • The metrics $g_{X,i}=h$ on
    $B^{h}_{2^{-i}}(p_{0})\cup\left(X\backslash B^{h}_{2^{i+1}}(p_{0})\right)\cup\bigsqcup\limits_{j=-i}^{i}B^{h}_{2^{j}+2^{-4i}}(p_{0})\backslash\bar{B}^{h}_{2^{j}-2^{-4i}}(p_{0})\subset(X,h),$
  • The Gromov-Hausdorff distance $d_{GH}\left((X\backslash\bar{\mathcal{W}}_{i}),d_{g_{X,i}}),(X,h)\right)\leq\left(1-2^{-i}\right)\epsilon$,
  • For any $p\in\{p_{0}\}\mathop{\cup}\limits_{j=0}^{\infty}\partial B^{h}_{2^{-j}}(p_{0})$, the Gromov-Hausdorff distance
    $r^{-1}d_{GH}\left((B^{h}_{r}(p)\backslash\bar{\mathcal{W}}_{i}),d_{g_{X,i}}),B^{\mathbb{R}^{n}}_{r}(0)\right)\leq\Psi(r|n,X,h),$
  • For any $x\in B^{h}_{2^{i}}(p_{0})\backslash\mathcal{W}_{i}\subset\left(X\backslash\mathcal{W}_{i},d_{g_{X,i}}\right)$, the distance $d_{g_{X,i}}(x,\mathcal{W}_{i})\leq 2^{2-i}$,
  • For any $i\in\mathbb{N}$, the identity mapping $(\partial W_{i},d_{g_{X,i+1}})\to(\partial W_{i},d_{g_{X,i}})$ is an $(1+2^{-i})$-bi-Lipschitz map,
  • For any $i,j\in\mathbb{N}$ and $x\in(X\backslash\bar{\mathcal{W}}_{i},d_{g_{X,i}})$, there exists a closed subset $K_{x,i,j}$ of $\mathbb{R}^{n}$ satisfies that $c_{j}^{-1}d_{GH}\left(B^{g_{X,i}}_{c_{j}}(x),K_{x,i,j}\right)=\Psi(j^{-1}|d_{h}(x,p_{0}),h)$,
  • For any positive integers $i\geq j$, $p\in\{p_{0}\}\bigsqcup\left(\bigsqcup\limits_{i=0}^{\infty}\partial B^{h}_{2^{-i}}(p_{0})\right)\bigsqcup\partial\mathcal{W}_{j}$, and $\sigma\in(0,1)$, the Gromov-Hausdorff distance
    $\sigma^{-1}d_{GH}(B_{\sigma}^{g_{X,i}}(p),B_{\sigma}^{g_{X,j}}(p))\leq 2^{-100j}.$
6. Proof of Theorem 1.1
Now we are ready to prove Theorem 1.1 and Corollary 1.1. For convenience, we restate Theorem 1.1 here.
Theorem 6.1.
Let $(X,h)$ be an $n$-dimensional smooth complete Riemannian manifold with $n\geq 3$ and $\mathrm{Ric}_{h}>\lambda\in\mathbb{R}$. Fix $\epsilon>0$. Then the sequence of $(n+2)$-dimensional manifolds $(M_{i},g_{i})\stackrel{{\scriptstyle GH}}{{\longrightarrow}}(X_{\epsilon},d_{\epsilon})$ we constructed in Proposition 5.1 satisfies the following properties.
  1. (a).
    $d_{GH}(X,X_{\epsilon})\leq\epsilon$,
  2. (b).
    The Ricci curvature $\mathrm{Ric}_{g_{i}}>\lambda$,
  3. (c).
    $(X_{\epsilon},d_{\epsilon})$ is $n$-rectifiable,
  4. (d).
    For any open subset $U$ of $X_{\epsilon}$, the homology group $H_{n-1}(U;\mathbb{Z})$ is infinitely generated. Consequently, every open set of $X_{\epsilon}$ is not a topological manifold.
Proof.
By the construction, one can easy to see that the sequence of $(n+2)$-dimensional manifolds $(M_{i},g_{i})\stackrel{{\scriptstyle GH}}{{\longrightarrow}}(X_{\epsilon},d_{\epsilon})$ satisfies (a) and (b).
By Proposition 5.1, for any $x\in X_{\epsilon}$, there exist a sequence of constants $c_{k}\to\infty$ and a sequence of closed subsets of $\mathbb{R}^{n}$, the Gromov-Hausdorff distance
$\lim_{k\to\infty}c_{k}^{-1}d_{GH}\left(B^{g_{X,i}}_{c_{k}}(x),K_{x,k}\right)=0.$
It follows that the rectifiable dimension of $X_{\epsilon}\leq n$ [6].
Since
$r^{-1}d_{GH}\left((B^{h}_{r}(p_{0})\backslash\bar{\mathcal{W}}_{i}),d_{g_{X,i}}),B^{\mathbb{R}^{n}}_{r}(0)\right)\leq\Psi(r|n,X,h),$
we see that $p_{0}\in\mathcal{R}_{n}(X_{\epsilon})$, and hence $R_{n}(X_{\epsilon})\neq\emptyset$. Thus the rectifiable dimension of $X_{\epsilon}\geq n$. Hence $(X_{\epsilon},d_{\epsilon})$ is $n$-rectifiable. This proves the property (c).
Now we consider the property (d). Let $U$ be an open subset of $X_{\epsilon}$. Assume that the homology group $H_{n}(U)$ is finitely generated. Without loss of generality, we can assume that $\mathrm{rank}(H_{n-1}(U))=N_{U}\in\mathbb{N}$.
Let $p_{\infty}\in U$ and $r_{\infty}>0$ such that $B_{2r_{\infty}}(p_{\infty})\subset U$. Choosing a sequence of points $p_{i}\in M_{i}$ converging to $p_{\infty}$. Then for any $N\in\mathbb{N}$, we can find $i_{N}\in\mathbb{N}$ such that for any $i\geq i_{N}$, the set $\mathcal{W}_{i}$ has at least $N$ connected components contained in $B_{2r_{\infty}}(p_{i})\subset U$. Denote them by $\mathcal{U}_{j}$, $j=1,\cdots,N$.
Now we recall that the inclusion map $I_{i,i-1}:X\backslash\bar{\mathcal{W}}_{i}\to X\backslash\bar{\mathcal{W}}_{i-1}$ is an $(1+2^{-i}\epsilon)$-Lipschitz map, and induces an $(1+2^{-i}\epsilon)$-bi-Lipschitz map on the connected components $\partial\mathcal{U}_{j}\to\partial\mathcal{U}_{j}$ of the boundary of $\partial\mathcal{W}_{i}$. Then we can see that the inclusion maps $I_{i_{N}+1,i_{N}}\circ\cdots\circ I_{i-1,i-2}\circ I_{i,i-1}$ converging to a Lipschitz map $X_{\epsilon}\to X\backslash\mathcal{W}_{i_{N}}$, and induces a bi-Lipschitz map between $\partial\mathcal{W}_{i_{N}}$ and a compact subset $\mathcal{Y}_{i_{N}}^{\infty}$ of $X_{\epsilon}$. Hence for $j=1,\cdots,N$, the image of $H_{n-1}(U)\to H_{n-1}(X\backslash\mathcal{W}_{i_{N}})$ induced by $U\subset X_{\epsilon}\to X\backslash\mathcal{W}_{i_{N}}$ contains the image of $\bigoplus_{j=1}^{N}H_{n-1}(\partial\mathcal{U}_{j})=H_{n-1}(\mathcal{W}_{i_{N}})\to H_{n-1}(X\backslash\mathcal{W}_{i_{N}})$ induced by the inclusion map $\bigsqcup_{j=1}^{N}\partial\mathcal{U}_{j}=\mathcal{W}_{i_{N}}\to X\backslash\mathcal{W}_{i_{N}}$. On the other hand, we have
$\mathrm{rank}\left(\mathrm{image}(H_{n-1}(U)\to H_{n-1}(X\backslash\mathcal{W}_{i_{N}}))\right)\leq\mathrm{rank}(H_{n-1}(U))\leq N_{U}.$
Hence we have
$\mathrm{rank}\left(\mathrm{image}\left(\bigoplus_{j=1}^{N}H_{n-1}(\partial\mathcal{U}_{j})\to H_{n-1}(X\backslash\mathcal{W}_{i_{N}})\right)\right)\leq\mathrm{rank}\left(\mathrm{image}(H_{n-1}(U)\to H_{n-1}(X\backslash\mathcal{W}_{i_{N}}))\right)\leq N_{U}.$
[…]
$\mathrm{rank}\left(\mathrm{image}\left(\bigoplus_{j=1}^{N}H_{n-1}(\partial\mathcal{U}_{j})\to H_{n-1}((X/\thicksim)\backslash\mathcal{W}_{i_{N}})\right)\right)\leq N_{U}.$
By the Mayer-Vietoris sequence [9, Section 2.2] given by $\cup_{j=1}^{N}\bar{\mathcal{U}}_{j}$ and $(X/\thicksim)\backslash\mathcal{W}_{i_{N}}$, we can get the following exact sequence:
$H_{n}(X/\thicksim)\to\bigoplus_{j=1}^{N}H_{n-1}(\partial\mathcal{U}_{j})\to H_{n-1}((X/\thicksim)\backslash\mathcal{W}_{i_{N}})\bigoplus H_{n-1}\left(\cup_{j=1}^{N}\bar{\mathcal{U}}_{j}\right)\to H_{n-1}(X/\thicksim).$
It follows that
$N=\mathrm{rank}\left(\bigoplus_{j=1}^{N}H_{n-1}(\partial\mathcal{U}_{j})\right)$$\leq$ $\mathrm{rank}\left(\mathrm{image}\left(\bigoplus_{j=1}^{N}H_{n-1}(\partial\mathcal{U}_{j})\to H_{n-1}((X/\thicksim)\backslash\mathcal{W}_{i_{N}})\right)\right)+\mathrm{rank}H_{n}(X/\thicksim)$
$\leq$ $N_{U}+\mathrm{rank}H_{n}(X/\thicksim)$
But obviously an arbitrarily large $N$ can be taken here, a contradiction. This proves (d), and the theorem follows. ∎
Then we prove Corollary 1.1.
Proof of Corollary 1.1: Let $(M_{i},g_{i})\stackrel{{\scriptstyle GH}}{{\longrightarrow}}(X_{\epsilon},d_{\epsilon})$ be the sequence of $(n+2)$-dimensional manifolds we constructed in Proposition 5.1. Then Proposition 5.1 implies that for any $x\in\mathcal{Y}_{j}^{\infty}$, all tangent cone at $x$ are isometric to $\mathbb{R}^{n-1}\times\mathbb{R}_{\geq 0}$, where $\mathcal{Y}_{j}^{\infty}$ is the limit of $\mathcal{W}_{j}$ in $X_{\epsilon}$. Hence $\cup_{j=1}^{\infty}\mathcal{Y}_{j}^{\infty}\subset\partial X_{\epsilon}$ is dense in $X_{\epsilon}$. Applying Proposition 5.1 again, one can conclude that $\cup_{i=1}^{\infty}\partial B^{h}_{2^{-i}}\subset R_{n}(p_{0})$. Then $\cup_{i=1}^{\infty}(\partial B^{h}_{2^{-i}}\cap\partial X_{\epsilon})=\emptyset$, and hence $\partial X_{\epsilon}$ has an infinite number of connected components. Thus $\partial_{\rm DPG}X_{\epsilon}=X_{\epsilon}$. Similarly, $\partial_{\rm KM}X_{\epsilon}$ has an infinite number of connected components. ∎